# An existence and uniqueness result to evolution equations with sign-changing pseudo-differential operators and its applications to logarithmic Laplacian operators and second-order differential operators without ellipticity

### v1

@@ Abstract.
This research was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No.2020R1A2C1A01003959)
@@ 1. Introduction
within a specific framework becomes highly challenging without imposing very strong assumptions on both the symbol $\sigma(x,\xi)$ and the data $f$. However, if one considers pseudo-differential operators which are independent of the space variable $x$, there might be intriguing theories for addressing a general symbol $\sigma(\xi)$ and data $f$ within a weak formulation. Specifically, when considering a polynomial symbol $\sigma(\xi)$, there exist exceptionally elegant theories regarding the existence and uniqueness of solutions to (elliptic) equations like (1.1) in the weak sense. Remarkably, these theories hold even in the absence of strong mathematical conditions on $\sigma$ and $f$. In particular, a solution $u$ exists in the sense of distributions if the symbol $\sigma$ is a polynomial, $\sigma(\xi)\not\equiv 0$, and $f$ is a distribution with a compact support on $\mathbb{R}^{d}$. Even more surprisingly, there always exists a distribution solution $u$ to (1.1) for any distribution $f$ if the symbol $\sigma$ is a non-zero polynomial, which is shown based on the Paley-Wiener theorem and duality arguments. This existence result holds even though a nonzero polynomial symbol $\sigma(\xi)$ does not satisfy an ellipticity condition, since the set of all zeros of a non-zero polynomial has a lower dimension. However, the uniqueness of a weak solution $u$ is not guaranteed if there is no ellipticity condition. Thus ellipticity is necessary and sufficient to guarantee both existence and uniqueness of a weak solution to (1.1) even for a polynomial symbol which is independent of $x$. Moreover, this well-posedness theory can be generalized to accommodate non-polynomial symbols $\sigma$ and tempered distributions $f$ by imposing supplementary conditions involving growth, regularity, and ellipticity on the symbol $\sigma$. For more in-depth information, we recommend that the reader consult the books [10, 11, 24, 25, 27]. It is noteworthy that for a polynomial symbol $\sigma(\xi)$, the sufficient condition to ensure the weak existence of (1.1) with $f$ having a distribution with compact support is simply that $\sigma(\xi)$ is not identically zero. This condition is feasible through a translation of the polynomial $\sigma(\xi)$ and using properties of convolution and analytic functions (see, for instance, [10, Theorem 7.3.10] and [11, Theorem 10.3.1]). However, this approach may not be applicable for non-polynomial symbols, particularly in scenarios where the symbols lack regularity. In such scenarios, an ellipticity condition on $\sigma$ becomes crucial, as it prompts the consideration of the reciprocal $\frac{1}{\sigma}$ to derive a fundamental solution (Green’s function).
@@
Drawing upon the criteria outlined in (1.7) and (1.8), extensive researches conducted by the authors and collaborators have consistently shown robust well-posedness for the equation (1.5). Examples of such research can be found in [1, 2, 5, 15, 16, 17, 18]. These operators exhibit a close connection to non-local operators and serve as generators of stochastic processes. Additionally, for a discussion on the strong well-posedness of these equations with non-local operators and generators of stochastic processes in $L_{p}$-spaces, we direct readers to the references [3, 4, 6, 12, 13, 14, 20, 22, 29, 30, 32].
@@
  • Let $\mathbb{N}$, $\mathbb{Z}$, $\mathbb{R}$, $\mathbb{C}$ denote the natural number system, the integer number system, the real number system, and the complex number system, respectively. For $d\in\mathbb{N}$, $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space.
  • For $i=1,...,d$, a multi-index $\alpha=(\alpha_{1},...,\alpha_{d})$ with $\alpha_{i}\in\{0,1,2,...\}$, and a function $g$, we set
@@
  • We employ the identical symbol “$\sup$” to represent both the supremum and the essential supremum by slightly abusing the notation $\sup$. On occasion, “$\operatorname*{ess\,sup}$” may be utilized to emphasize its role as the essential supremum.
  • We use $\mathcal{D}(\mathbb{R}^{d})$ to denote the space of all infinitely differentiable functions with compact supports. $\mathcal{S}(\mathbb{R}^{d})$ represents the Schwartz space on $\mathbb{R}^{d}$ and the topology on $\mathcal{S}(\mathbb{R}^{d})$ is generated by the Schwartz semi-norms $\sup_{x\in\mathbb{R}^{d}}|x^{\alpha}(D^{\beta}f)(x)|$ for all multi-indexes $\alpha$ and $\beta$. $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is used to denote the dual space of $\mathcal{S}(\mathbb{R}^{d})$, i.e. $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is the space of all tempered distributions on $\mathbb{R}^{d}$. Additionally, we assume that $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is a topological space equipped with the weak*-topology if there is no special remark about the topology on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$.
@@
[…]
@@
    For the sake of simplicity, we often omit the subscripts denoted by $x$ and $\xi$. It is well-known that both the Fourier transform and the inverse Fourier transform have extensions on $L_{1}(\mathbb{R}^{d})+L_{2}(\mathbb{R}^{d})$. By slightly abusing the notation, we use the same notation $\mathcal{F}$ and $\mathcal{F}^{-1}$ to denote these extensions. We discuss more details some properties of these extensions and corresponding inversion theorems in Section 3.
@@
  • For $z\in\mathbb{C}$, $\Re[z]$ denotes the real part of $z$, $\Im[z]$ is the imaginary part of $z$ and $\bar{z}$ is the complex conjugate of $z$.
@@ 2. Settings and main results
[…]
@@
Our main goal of this paper is to show the existence and uniqueness of a weak solution $u$ to the Cauchy problem (1.5). Note that it is not easy to recover a strong solution to (1.5) even in $L_{p}$-spaces if there are no strong mathematical conditions on the symbol and data. A naive condition is that for each $t\in(0,T)$, $\psi(t,\cdot)\mathcal{F}[u](t,\cdot)\in L_{p}(\mathbb{R}^{d})$ with $p\in[1,2]$. Then the function $t\mapsto\psi(t,-i\nabla)u(t,\cdot)$ is well-defined as an $L_{p^{\prime}}(\mathbb{R}^{d})$-valued function defined on $(0,T)$ due to the Riesz-Thorin Theorem, where $p^{\prime}$ denotes the Hölder conjugate of $p$, i.e.
@@
[…]
@@
[…]
@@
Since $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\subset\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, there is the subspace topology on $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and the Borel sets generated by this topology. Additionally, this identification with a locally integrable function could be easily performed in any domain of $\mathbb{R}^{d}$ due to the localizations of distributions. More precisely, if there exist a open set $\mathcal{U}\subset\mathbb{R}^{d}$ and a locally integrable function $v$ on $\mathcal{U}$ so that
@@
then we identify $\mathcal{F}[u]$ with $v$ on $\mathcal{U}$, where $\mathcal{D}\left(\mathcal{U}\right)$ denotes the set of all infinitely differential functions on $\mathbb{R}^{d}$ with compact supports in $\mathcal{U}$. In such a case, we say that $\mathcal{F}[u]$ has a realization $v$ on $\mathcal{U}$ and consider $\mathcal{F}[u]$ as a locally integrable function on $\mathcal{U}$. In other words, for any $u$ in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, we say that $\mathcal{F}[u]$ has a realization $v$ on $\mathcal{U}$ or simply $\mathcal{F}[u]$ is locally integrable on $\mathcal{U}$ if (2.3) holds.
@@
[…]
@@
[…]
@@
[…]
@@
where $\operatorname{supp}\psi(s,\cdot)=\overline{\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$, $\overline{\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$ is the closure of ${\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$. However, we can give any value to $\mathcal{F}[u(s,\cdot)]$ outside of $\operatorname{supp}\psi(s,\cdot)$ since it does not influence on the terms in (2.6). Thus it might suggest that the space $L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ is not appropriate to guarantee the uniqueness of a solution $u$. By the way, there is a typical action to extend by putting
@@
[…]
@@
We now turn our attention to a significant subset within $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, which comprises functions $u$ whose Fourier transform is in an weighted local $L_{p,q}$-space. This subset can be formally expressed as:
@@
There are numerous mathematical conveniences when we exclusively consider positive weights. Nevertheless, it is crucial not to disregard the possibility of weights becoming zero, as this allows us to include the interesting cases in which the symbol $\psi(t,\xi)$ may become degenerate. Furthermore, it is true that, for all values of $p$ and $q$ within the range of $[1,\infty]$,
@@
[…]
@@
We can describe a function $u$ belonging to the space $L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ as having a strong local integrability with respect to the variable $t$ and a weaker form of local integrability with respect to the variable $x$. To clarify further, consider a topological space denoted by $X$, and recall that when we say a property holds locally, it means that the property is satisfied for all compact subsets $K\subset X$. In this context, our function $u\in L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ can be described as locally integrable with respect to $x$, but it satisfies a somewhat stronger condition compared to the conventional notion of local integrability concerning the time variable $t$. This distinction is why we employed different notations, namely, “$loc$” and ‘$\ell oc$”, to indicate the local integrability properties of $t$ and $x$, respectively.
@@
$\int_{B_{R}}\exp\left(\int_{0}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi+\int_{B_{R}}\int_{0}^{t}\exp\left(\int_{s}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi$ <<$\infty$,< ∞ , (2.9)
@@
The solution $u$ to (1.5) in Theorem 2.21 is not defined as a complex-valued function on $[0,T)\times\mathbb{R}^{d}$. It is a function in the class $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, which is clarified by the action in (2.11). Therefore, formally, we may say that
@@
This kernel representation shows that it is not hopeful to expect that $u$ has a realization on $[0,T)\times\mathbb{R}^{d}$ if the symbol $\psi(t,\xi)$ is not negative.
@@
where $i$ beside $b^{j}(t)$ is the imaginary number so that $i^{2}=-1$. To maintain clarity and facilitate readers’ comprehension, we provide a modified version of the definition of a Fourier-space weak solution for second-order cases. This indicates that the structure of our weak formulation closely parallels that of the classical one.
@@
Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in

### v2

@@ Abstract.
J.-H. Choi has been supported by a KIAS Individual Grant (MG102701) at Korea Institute for Advanced Study. I. Kim has been supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No.2020R1A2C1A01003959)
@@ 1. Introduction
within a specific framework becomes highly challenging without imposing very strong assumptions on both the symbol $\sigma(x,\xi)$ and the data $f$. However, if one considers pseudo-differential operators which are independent of the space variable $x$, there might be intriguing theories for addressing a general symbol $\sigma(\xi)$ and data $f$ within a weak formulation. Specifically, when considering a polynomial symbol $\sigma(\xi)$, there exist exceptionally elegant theories regarding the existence and uniqueness of solutions to (elliptic) equations like (1.1) in the weak sense. Remarkably, these theories hold even in the absence of strong mathematical conditions on $\sigma$ and $f$. In particular, a solution $u$ exists in the sense of distributions if the symbol $\sigma$ is a polynomial, $\sigma(\xi)\not\equiv 0$, and $f$ is a distribution with a compact support in $\mathbb{R}^{d}$. Even more surprisingly, there always exists a distribution solution $u$ to (1.1) for any distribution $f$ if the symbol $\sigma$ is a non-zero polynomial, which is shown based on the Paley-Wiener theorem and duality arguments. This existence result holds even though a nonzero polynomial symbol $\sigma(\xi)$ does not satisfy an ellipticity condition, since the set of all zeros of a non-zero polynomial has a lower dimension. However, the uniqueness of a weak solution $u$ is not guaranteed if there is no ellipticity condition. Thus ellipticity is necessary and sufficient to guarantee both existence and uniqueness of a weak solution to (1.1) even for a polynomial symbol which is independent of $x$. Moreover, this well-posedness theory can be generalized to accommodate non-polynomial symbols $\sigma$ and tempered distributions $f$ by imposing supplementary conditions involving growth, regularity, and ellipticity on the symbol $\sigma$. For more in-depth information, we recommend that the reader consult the books [10, 11, 24, 25, 27]. It is noteworthy that for a polynomial symbol $\sigma(\xi)$, the sufficient condition to ensure the weak existence of (1.1) with $f$ having a distribution with compact support is simply that $\sigma(\xi)$ is not identically zero. This condition is feasible through a translation of the polynomial $\sigma(\xi)$ and using properties of convolution and analytic functions (see, for instance, [10, Theorem 7.3.10] and [11, Theorem 10.3.1]). However, this approach may not be applicable for non-polynomial symbols, particularly in scenarios where the symbols lack regularity. In such scenarios, an ellipticity condition on $\sigma$ becomes crucial, as it prompts the consideration of the reciprocal $\frac{1}{\sigma}$ to derive a fundamental solution (Green’s function).
@@
Drawing upon the criteria outlined in (1.7) and (1.8), extensive research conducted by the authors and collaborators have consistently shown robust well-posedness for the equation (1.5). Examples of such research can be found in [1, 2, 5, 15, 16, 17, 18]. These operators exhibit a close connection to non-local operators and serve as generators of stochastic processes. Additionally, for a discussion on the strong well-posedness of these equations with non-local operators and generators of stochastic processes in $L_{p}$-spaces, we direct readers to the references [3, 4, 6, 12, 13, 14, 20, 22, 29, 30, 32].
@@
  • Let $\mathbb{N}$, $\mathbb{Z}$, $\mathbb{R}$, $\mathbb{C}$ denote the natural number system, the integer number system, the real number system, and the complex number system, respectively. For $\in\mathbb{N}, d\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space.
  • For i=1,,di=1,...,d, a multi-index $\alpha=(\alpha_{1},...,\alpha_{d})$ with $\alpha_{i}\in\{0,1,2,$...\}, and a function $g$, we set
@@
  • We employ the identical symbol $\sup$” to represent both the supremum and the essential supremum by slightly abusing the notation $\sup$. On occasion, $\operatorname*{ess\,sup}$” may be utilized to emphasize its role as the essential supremum.
  • We use $\mathcal{D}(\mathbb{R}^{d})$ to denote the space of all infinitely differentiable functions with compact supports. $\mathcal{S}(\mathbb{R}^{d})$ represents the Schwartz space on $\mathbb{R}^{d}$ and the topology on $\mathcal{S}(\mathbb{R}^{d})$ is generated by the Schwartz semi-norms $\sup_{x\in\mathbb{R}^{d}}|x^{\alpha}(D^{\beta}f)(x)|$ for all multi-indices $\alpha$ and $\beta$. $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is used to denote the dual space of $\mathcal{S}(\mathbb{R}^{d})$, i.e. $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is the space of all tempered distributions on $\mathbb{R}^{d}$. Additionally, we assume that $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is a topological space equipped with the weak*-topology if there is no special remark about the topology on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$.
@@
    if and only if there exists a measurable subset $\mathcal{T}\subset\mathcal{U}$ such that the Lebesgue measure $|\mathcal{U}\setminus\mathcal{T}|$ is zero and $f(t)=g(t)$ for all $t\in\mathcal{T}$. Moreover, we say that a function $f(t)$ is defined $(a.e.)$ on a Lebesgue measurable (or Borel measurable) set $\mathcal{U}$ if there exists a measurable subset $\mathcal{T}\subset\mathcal{U}$ such that the Lebesgue measure $|\mathcal{U}\setminus\mathcal{T}|=0$ and $f(t)$ is defined for all $t\in\mathcal{T}$.
@@
    For the sake of simplicity, we often omit the subscripts denoted by $x$ and $\xi$. It is well-known that both the Fourier transform and the inverse Fourier transform have extensions on $L_{1}(\mathbb{R}^{d})+L_{2}(\mathbb{R}^{d})$. By slightly abusing the notation, we use the same notation $\mathcal{F}$ and $\mathcal{F}^{-1}$ to denote these extensions. We discuss further details and some properties of these extensions and corresponding inversion theorems in Section 3.
@@
  • For $\in\mathbb{C}, [z]\Re[z$] denotes the real part of $z$, $\Im[z]$ is the imaginary part of $z$ and $\bar{z}$ is the complex conjugate of $z$.
@@ 2. Settings and main results
We fix $T\in(0,\infty]$ and $d\in\mathbb{N}$ throughout the paper. Here $T$ and $d$ denote the terminal time of the evolution equation and the dimension of the space-variable, respectively. Note that $T=\infty$ is possible in our theory. All functions are complex-valued if there is no special remark about the range of a function. In particular, $\psi(t,\xi)$ denotes a complex-valued measurable function defined $(a.e.)$ on $(0,T)\times\mathbb{R}^{d}$ and denote
@@
The main goal of this paper is to show the existence and uniqueness of a weak solution $u$ to the Cauchy problem (1.5). Note that it is not easy to recover a strong solution to (1.5) even in $L_{p}$-spaces if there are no strong mathematical conditions on the symbol and data. A naive condition is that for each $t\in(0,T)$, $\psi(t,\cdot)\mathcal{F}[u](t,\cdot)\in L_{p}(\mathbb{R}^{d})$ with $p\in[1,2]$. Then the function $t\mapsto\psi(t,-i\nabla)u(t,\cdot)$ is well-defined as an $L_{p^{\prime}}(\mathbb{R}^{d})$-valued function defined on $(0,T)$ due to the Riesz-Thorin Theorem, where $p^{\prime}$ denotes the Hölder conjugate of $p$, i.e.
@@
then we identify $\mathcal{F}[u]=v$ and consider $\mathcal{F}[u]$ as a function defined $(a.e.)$ on $\mathbb{R}^{d}$, where $\overline{\varphi(\xi)}$ denotes the complex conjugate of $\varphi(\xi)$. In this case, we usually say that the distribution $\mathcal{F}[u]$ has the realization $v$ on $\mathbb{R}^{d}$. This identification even works for any element $u\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$. In other words, we say that $u\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ is locally integrable if and only if there is a $v\in L_{1,\ell oc}(\mathbb{R}^{d})$
@@
This function $v$ is uniquely determined $(a.e.)$ on $\mathbb{R}^{d}$ since both $\mathcal{D}(\mathbb{R}^{d})$ and $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ are dense in $L_{\infty}(B_{R})$ for all $R\in(0,\infty)$ (by considering restrictions of $\mathcal{D}(\mathbb{R}^{d})$ or $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ on each $B_{R}$). Due to this identification, we could introduce an important subclass of $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ consisting of all functions whose Fourier transform is locally integrable. Here is the precise definition. We use $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ to denote the subspace of $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ whose Fourier transform is in $L_{1,\ell oc}(\mathbb{R}^{d})$, i.e. $u\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ if and only if
@@
Since $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\subset\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, there is the subspace topology on $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and the Borel sets generated by this topology. Additionally, this identification with a locally integrable function could be easily performed in any domain of $\mathbb{R}^{d}$ due to the localizations of distributions. More precisely, if there exist an open set $\mathcal{U}\subset\mathbb{R}^{d}$ and a locally integrable function $v$ on $\mathcal{U}$ so that
@@
then we identify $\mathcal{F}[u]$ with $v$ on $\mathcal{U}$, where $\mathcal{D}\left(\mathcal{U}\right)$ denotes the set of all infinitely differentiable functions on $\mathbb{R}^{d}$ with compact supports in $\mathcal{U}$. In such a case, we say that $\mathcal{F}[u]$ has a realization $v$ on $\mathcal{U}$ and consider $\mathcal{F}[u]$ as a locally integrable function on $\mathcal{U}$. In other words, for any $u$ in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, we say that $\mathcal{F}[u]$ has a realization $v$ on $\mathcal{U}$ or simply $\mathcal{F}[u]$ is locally integrable on $\mathcal{U}$ if (2.3) holds.
@@
$L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ is used to denote the class of all Borel measurable functions $f$ with values in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ defined $(a.e.)$ on $(0,T)$ such that
@@
We define $L_{0}\left((0,T);\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\right)$ as the subspace of $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ consisting of all Borel measurable functions $f$ with values in $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ defined $(a.e.)$ on $(0,T)$ so that
@@
We write $f\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\right)$ if $f$ is a $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$-valued Borel measurable function defined $(a.e.)$ on $(0,T)$ such that
@@
where $\operatorname{supp}\psi(s,\cdot)=\overline{\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$, $\overline{\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$ is the closure of $\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}$}. However, we can give any value to $\mathcal{F}[u(s,\cdot)]$ outside of $\operatorname{supp}\psi(s,\cdot)$ since it does not affect the terms in (2.6). Thus it might suggest that the space $L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ is not appropriate to guarantee the uniqueness of a solution $u$. By the way, there is a typical action to extend by putting
@@
Let $p,q\in[1,\infty)$, $w$, and $W$ be non-negative (Borel) measurable functions defined $(a.e.)$ on $(0,T)$ and $(0,T)\times\mathbb{R}^{d}$, respectively. We write $u\in L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)$ if and only if
@@
We now turn our attention to a significant subset within $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, which comprises functions $u$ whose Fourier transform is in a weighted local $L_{p,q}$-space. This subset can be formally expressed as:
@@
There are numerous mathematical conveniences when we exclusively consider positive weights. Nevertheless, it is crucial not to disregard the possibility of weights becoming zero, as this allows us to include the interesting cases in which the symbol $\psi(t,\xi)$ may become degenerate. Furthermore, it is true that, for all values of $p$ and $q$ within the range of $1,\infty]$,
@@
for all $p,q\in[1,\infty]$ since $|\mathcal{F}[u(t,\cdot)](\xi)|<\infty$ $(a.e.)$ $(t,\xi)\in(0,T)\times\mathbb{R}^{d}$ based on realizations on $\mathbb{R}^{d}$ for any
@@
We can describe a function $u$ belonging to the space $L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ as having a strong local integrability with respect to the variable $t$ and a weaker form of local integrability with respect to the variable $x$. To clarify further, consider a topological space denoted by $X$, and recall that when we say a property holds locally, it means that the property is satisfied for all compact subsets $K\subset X$. In this context, our function $u\in L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ can be described as locally integrable with respect to $x$, but it satisfies a somewhat stronger condition compared to the conventional notion of local integrability concerning the time variable $t$. This distinction is why we employed different notations, namely, “locloc” and $\ell oc$”, to indicate the local integrability properties of $t$ and $x$, respectively.
@@
$\int_{B_{R}}\exp\left(\int_{0}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi+\int_{B_{R}}\int_{0}^{t}\exp\left(\int_{s}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi$ $<\infty,$ (2.9)
@@
The solution $u$ to (1.5) in Theorem 2.21 is not defined as a complex-valued function on $0,T)\times\mathbb{R}^{d}$. It is a function in the class $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, which is clarified by the action in (2.11). Therefore, formally, we may say that
@@
This kernel representation shows that it is not hopeful to expect that $u$ has a realization on $0,T)\times\mathbb{R}^{d}$ if the symbol $\psi(t,\xi)$ is not negative.
@@
where $i$ beside $b^{j}(t)$ is the imaginary unit so that $i^{2}=-1$. To maintain clarity and facilitate readers’ comprehension, we provide a modified version of the definition of a Fourier-space weak solution for second-order cases. This indicates that the structure of our weak formulation closely parallels that of the classical one.
@@
Then there exists a unique weak solution $u$ ((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in